OpenStudy (anonymous):

Sue has four $20 bills in her wallet. She needs $200 in cash to buy a new MP3 player. How many more $20 bills does Sue need to buy the MP3 player? Sue needs more $20 bills.

OpenStudy (anonymous):

Is this multiple choice? Or did you just answer it yourself?

OpenStudy (anonymous):

IT'S 10 RIGHT

OpenStudy (anonymous):

When they say bills are they referring yo dollar bills? or twenty?

OpenStudy (anonymous):

IT TEN BECAUSE 10 *20 =200 CORRECT

OpenStudy (anonymous):

Not exactly....Give me a sec

OpenStudy (anonymous):

IT'S IS 10

OpenStudy (anonymous):

So if you have 4 20 dollar bills. So 20+20+20+20 which is 80. So she has 80 dollars right now but she needs 200

OpenStudy (anonymous):

We can subtract 80 from 200 to get 120. The last step is to divide 120 by 20 since it needs to be 20 bills and you get 6

OpenStudy (anonymous):

Sue need 6 more twenty bills to buy the mp3 player

OpenStudy (anonymous):

YES SO 20+20+20+20+20+20+20+20+20+20=200

OpenStudy (anonymous):

Yes that how much twenty's she needs. But is says HOW MANY MORE DOES SHE NEED.

OpenStudy (anonymous):

So what you did was right.. 10 twentys but she already has four. So 10-4=6

OpenStudy (anonymous):

OH YOUR RIGHT

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

Ok glad to help :)